### v1

@@ 4.3 Mass and Chemical Disorder
Figure 8: Bonding environment for Si atoms with (a) the highest and (b) the lowest SPTC in the Si0.75Ge0.25 crystal. In (a), the silicon atom (1) has the highest SPTC and is connected to silicon atoms (2, 3, 4, 5), which are also connected to other silicon atoms. In (b), the silicon atom (1), which has the lowest SPTC among silicon atoms in the system, is connected to two silicon atoms (2, 3) and two germanium atoms (4, 5).
Further, atomistic analysis of the Si0.75Ge0.25 crystal reveals distinct correlations between SPTC and local environment of individual atoms. For example, a silicon atom exhibits the highest SPTC when connected to four neighboring silicons, which in turn are also connected with other silicon atoms. Conversely, a silicon atom exhibits the lowest SPTC when bonded to the germanium atoms and/or connected to silicon atoms which are connected to germanium atoms. Figure 8, provides a visual description of thermal characteristics of silicon atoms in different bonding environments within the Si-Ge system.
@@ 5 Conclusion
This study introduces the Space-Projected Thermal Conductivity (SPTC) methodology that employs Green-Kubo theory to map TC onto atomic sites, thus identifying more or less thermally active regions. To illustrate the utility of this approach, we calculated the SPTC to carbon, silicon, and silicon-germanium systems, showing that point, planar (grain boundary) and mass defects in these systems considerably reduce local TC. Although the developed methodology does not account for electronic contributions, limiting quantitative comparisons with experimental data, it effectively predicts thermal conduction paths consistent with known defect behavior in these materials.
Funding
This project was supported by ONR Grant N00014-23-1-2773. This work also used computational resources from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, through allocation DMR190002, MAT240030, funded by the U.S. National Science Foundation (NSF) grant number: 2138259, 2138286, 2138307, 2137603, and 2138296.

### v2

@@ 4.3 Mass and Chemical Disorder
Further, atomistic analysis of the Si0.75Ge0.25 crystal reveals distinct correlations between SPTC and the local environment of individual atoms. For example, a silicon atom exhibits the highest SPTC when connected to four neighboring silicons, which are also connected with other silicon atoms. Conversely, a silicon atom exhibits the lowest SPTC when bonded to the germanium atoms and/or connected to silicon atoms which are connected to germanium atoms. Figure 8, provides a visual description of the thermal characteristics of silicon atoms in different bonding environments within the Si-Ge system.
@@ 5 Conclusion
This study introduces the Site-Projected Thermal Conductivity (SPTC) methodology that employs the Green-Kubo theory to map TC onto atomic sites, thus identifying more or less thermally active regions. To illustrate the utility of this approach, we calculated the SPTC to carbon, silicon, and silicon-germanium systems, showing that point, planar (grain boundary), and mass defects in these systems considerably reduce local TC. Although the developed methodology does not account for electronic contributions, limiting quantitative comparisons with experimental data, it effectively predicts thermal conduction paths consistent with known defect behavior in these materials. To date, we have implemented the method “exactly”, However, since the dynamical matrix is short-ranged [32], it may become possible to explore much larger systems.
Acknowledgment
The authors thank Prof. J. J. Dong for helpful discussions. This research is supported by ONR Grant N00014-23-1-2773. It also used computational resources from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, through allocation DMR190002, MAT240030, funded by the U.S. National Science Foundation (NSF) grant number: 2138259, 2138286, 2138307, 2137603, and 2138296.